OpenStudy (anonymous):

You research the average cost of whole milk for several recent years to look for trends. The table shows your data. What is the equation for a line of best fit? How much would you expect to pay for a gallon of whole milk in the year 2023 the chart is in the comments

OpenStudy (dumbcow):

is this a stats class?

OpenStudy (anonymous):

no algabra 2

OpenStudy (dumbcow):

oh ok, that makes the problem a bit easier from the table data, select smallest and largest costs use those as 2 points to get equation of a line

OpenStudy (anonymous):

can u type that out?

OpenStudy (dumbcow):

or wait...hmm have you done regression with your graphing calculators in class ?

OpenStudy (anonymous):

no.... it has to do with slope intercept form... im just not sure how to get the equation

OpenStudy (anonymous):

this is the whole question 5. You research the average cost of whole milk for several recent years to look for trends. The table shows your data. What is the equation for a line of best fit? How much would you expect to pay for a gallon of whole milk in the year 2023? (1 point)y = 2.54x + 0.098; $4.98 y = 0.098x + 2.54; $6.16 y = 2054x + 0.098; $6.16 y = 0.098x + 2.54; $4.98

OpenStudy (dumbcow):

ok sorry not sure because it says "best fit" slope intercept form is : y = mx + b all you need is 2 points use 1st and last data points --> (1998 , 2.54) and (2008, 3.72)

OpenStudy (anonymous):

what would i do with those poiunts thats where i get stuck

OpenStudy (anonymous):

points*

OpenStudy (dumbcow):

whenever they give years, change the x_values, set initial point as first year so "x" represents number of years after 1998 1998 --> x=0 2008 --> x=10 (0,2.54) and (10,3.72) find slope: \[m = \frac{y_2 -y_1}{x_2 - x_1} = \frac{3.72 - 2.54}{10-0} = \frac{1.18}{10} = .118\] y-intercept is always "y" value when x=0 b = 2.54 --> y = .118x +2.54

OpenStudy (dumbcow):

the regression line is: y = .102x +2.5 i don't know how they got a slope .098 ??

OpenStudy (dumbcow):

since this is easy multiple choice though....answer is clearly option D

OpenStudy (anonymous):

It wasnt D it was C. You had it right the first time.